OpenStudy (anonymous):

In right triangle PQR where R is the right angle, PQ=13, QR=12, what is the length of PR ?

OpenStudy (anonymous):

a^2+b^2=c^2 that is the equation for finding one length when you know two others. in the formula, C is the hypotenuse. since we know that R is the right angle, and the hypotenuse is always opposite the right angle, then we know that PQ is the hypotenuse. so we get: a^2 + 12^2 = 13^2 now we solve a^2= 13^2 - 12^2 a^2 = 169 - 144 a^2 = 25 a=5 so the length of PR is 5

OpenStudy (anonymous):

wow. thanks for the help(: